# LOTUS: Enabling Semantic Queries with LLMs Over Tables of Unstructured and Structured Data

### v1

@@ 2.2. Semantic Operators
LOTUS also exposes advanced relevance-based re-ranking functionality for search. Users can specify a re-ranker model, as shown in Figure 15, and then set the n_rerank parameter during the semantic search. The semantic search in this case will first find the top- most relevant documents according to the retriever model, and then re-rank the top- found documents to return the top n_rerank ones using the re-ranker model.
Sem_cluster_by assigns a group to each row of the table according to semantic similarity. This operator is akin to a relational group_by, but uses semantic similarity to determine the groups, rather than equality over the specified column. As Figure 16 shows, the user specifies an indexed column to cluster, and the number of clusters to create. The returned table augments the input table with an additional column that specifies the assigned cluster_id of each row. Users can optionally also return the similarity scores between each row and its cluster centroid in an additional column of the returned table.
3. Implementing and Optimizing Semantic Operators
Semantic operators create a rich design space, allowing for a diverse set of algorithmic decisions and optimizations which have significant consequences on system efficiency and accuracy. We identify multiple possible algorithms for these operators, noting their performance implications. LOTUS’ query engine aims to automatically exploit this rich design space to provide an optimized implementation of each individual operator and composite pipelines.
Specifically, LOTUS employs a series of novel algorithms designed to leverage the semantics of each operator in order to maximize parallel batched-inference opportunities, use model cascades (Viola and Jones, 2001; Yue et al., 2024; Chen et al., 2023; Kang et al., 2017, 2022) with a lightweight scoring function unique to our setting, leverage the semantic similarity index, and perform algorithmic approximations for expensive operations (e.g. joins). We see in Section 4 that these decisions can significantly improve efficiency while maintaining high quality results for LOTUS’ query pipelines.
@@ 3. Implementing and Optimizing Semantic Operators
While our initial investigation studies several operator-specific design decisions unique to the semantic bulk processing setting, we envision a rich set of additional optimization opportunities. Several prior works demonstrate performance gains in both logical query plan optimizations (e.g. operator re-ordering (Liu et al., 2024c, b; Lin et al., 2024; Lu et al., 2018)) and other general LM-approximation techniques (e.g. code synthesis (Arora et al., 2023; Liu et al., 2024b) and prompt adaptation (Chen et al., 2023)), which we find to be promising opportunities for LOTUS’ implementation, left to future work.
LOTUS’ current implementation extends the Pandas API (pan, [n. d.]). We leverage vLLM (Kwon et al., 2023) to perform efficient batched inference, and we use FAISS (Johnson et al., 2017) for efficient vector similarity search.
@@
LOTUS’ semantic filter runs batched LLM calls over a set of rows, prompting the model to output a boolean value for each row to indicate whether the record passes the user’s natural language predicate. Figure 18 shows an example of the operator-specific filtering instruction, which is managed transparently by the system.
@@ 3.1.1. Optimizations
We leverage model cascades, using an efficient scoring function unique to our setting, to further optimize the filter function. Many prior works have studied model cascades for ML tasks (Viola and Jones, 2001; Yue et al., 2024; Chen et al., 2023; Kang et al., 2017, 2022), leveraging a relatively cheap and inaccurate proxy model, along with an accurate but expensive oracle model in order to reduce the inference cost by routing easy queries to the small model, and resorting to the large model where necessary. Typically, a scoring function provides confidence scores over the proxy model’s outputs, and allows the system to decide when to route examples to the larger model. Applying this paradigm to LMs introduces several challenges, including generating a reliable scoring function. Several works study this problem, suggesting to fine-tune a smaller LM to score each question along with the answer produced by the weak LLM (Chen et al., 2023), or invoking multiple sampling paths from the small LM and evaluating self-consistency to build a confidence score (Yue et al., 2024). While these methods provide general-purpose scoring functions, LOTUS’ scoring function leverages the operator semantics to provide a lighter-weight scoring function, which avoids the execution time penalty of scoring with an additional model or re-sampling from the proxy-model.
Specifically, LOTUS’ cascade scoring function leverages the binary-label output of each LM invocation. The system first batch-processes each record using the cheaper LM, then computes confidence scores by exponentiation the log-probabilities for the output LLM tokens corresponding to the True or False answer.
For records with confidence scores below the user-defined threshold, the system then passes these records in batch to the larger LM. As we show, in Section 4, leveraging model cascades in this way for semantic filtering can significantly improves the system’s efficiency while maintaining high quality results for some tasks.
3.2. sem_topk
Performing a semantic top- ranking requires logically reasoning across rows, entailing joint reasoning over often large amounts of data. Implementing an efficient algorithm introduces many important design decisions pertaining to managing LM context length limits, grouping documents within each LM context window to achieve high quality LM-based comparisons over subsets of the data, and choosing an efficient top-k ranking algorithm to combine LM-based comparisons. These decisions have notable performance implications on efficiency and result quality, as we show in Section 4. While prior works have studied LM-based passage re-ranking (Desai and Durrett, 2020; Drozdov et al., 2023; Liang et al., 2022; Ma et al., 2023; Pradeep et al., 2023a, b; Qin et al., 2024; Sachan et al., 2022; Sun et al., 2023) and ranking with noisy comparisons (Shah and Wainwright, 2016; Braverman and Mossel, [n. d.]) with the goal of achieving high quality results in a modest number of total LM calls or comparisons, LOTUS’ implementations provides a generalized ranking algorithms for arbitrary natural language expressions and aims to produce both high quality results and low query execution time. We leverage several algorithmic design decisions and optimizations to achieve this.
3.2.1. Algorithms
Prior LM-based sorting and ranking works suggest several methods for performing LM-based comparisons. Specifically, prior works primarily study three classes of methods: point-wise ranking methods (Desai and Durrett, 2020; Drozdov et al., 2023; Liang et al., 2022; Sachan et al., 2022; Wu et al., 2024), list-wise ranking methods (Ma et al., 2023; Pradeep et al., 2023a, b; Sun et al., 2023), and pair-wise ranking methods (Shah and Wainwright, 2016; Qin et al., 2024). Point-wise methods output a relevance score for each document independently; while these methods are fully parallelizable across rows, prior work (Desai and Durrett, 2020) shows they provide poor accuracy due to difficulty in calibrating scores across prompts. We note that these methods can be implemented in LOTUS with a sem_map, but are not the focus of our semantic top- operator due to their quality limitations. By contrast, list-wise methods feed a partial list of 10 to 20 documents to the LM and prompt it to output a ranking over these, which can be aggregated using a sliding window approach. Unfortunately, prior works shows these methods are often prone to prediction failures (Sun et al., 2023) and sensitive to document ordering in the prompt (Qin et al., 2024).
Pairwise-prompting methods instead offer a simple and effective approach that feeds a single pair of documents to the LM in each invocation, prompting the model to perform a comparison and output a binary label. This method has been shown to be an effective base unit for ranking with relatively high robustness to input order sensitivity (Qin et al., 2024).
To aggregate pairwise comparisons, we consider several ranking algorithms, including a quadratic sorting algorithm, a heap-based top-k algorithm and a quick-select-based top-k ranking algorithm. The quadratic ranking algorithm obtains a comparison between each pair of input documents, and uses a win rate to determine a ranking over all elements before selecting the best. In contrast, the heap-based algorithm maintains a heap of size , and makes a linear pass over the data, updating the heap when more promising elements are found. Each time a new element is inserted or removed from the heap, the algorithm performs a series of sequential LM comparisons to update the data structure. Lastly the quick-select based algorithm proceeds in successive rounds, each time choosing a pivot, and comparing all other remaining elements in the document set to the pivot item to determine the rank of the pivot. Because each round is fully parallelizable, we perform these LM-based comparisons efficiently in batch before recursing in the next round.
3.2.2. Discussion
As the core LM-based computation unit over subsets of the data, LOTUS’ current implementation uses pair-wise rankings, following recent prior work (Qin et al., 2024), which demonstrates its effectiveness. We also find this choice useful for implementing further optimizations, such as model cascades, described below. Figure 19 shows an example task-instruction prompt, which we use to implement the LM-based comparison by instructing the model to select between two documents according to the user-defined sorting criteria.
Additionally, for aggregating comparisons, LOTUS’ current implementation leverages the quick-select-based (Hoare, 1961) top-k ranking algorithm, by default. Prior work (Qin et al., 2024) studies several alternative algorithms, including the heap-based ranking implementation and the quadratic sorting algorithm. We study these alternatives in Section 4, and find that the quick-select-based algorithm offers high accuracy while also offering an efficient implementation with at least an order magnitude fewer total calls then the quadratic sorting algorithm and more opportunities for batched calls leading to lower execution time compared to a heap-based implementation. We believe future work may leverage these multiple algorithms to allow the user to declaratively trade-off between performance metrics, like accuracy, execution time, and cost.
3.2.3. Optimizations
LOTUS’ top- algorithm is amenable to several optimizations. First, the top- algorithm’s pair-wise comparisons each invoke the LM to output a binary label, which is amenable to model cascades. To implement this, we use a simple an efficient scoring procedure based on log-probabilities of the generated tokens. We describe this procedure above for sem_filter (Section 3.1.1) and apply an equivalent procedure here. In Section 4, we demonstrate this optimization allows programmers to leverage small, cheap models in conjunction with large, expensive models to obtain high accuracy results at reduced cost.
@@ 3.2.3. Optimizations
Additionally, LOTUS can leverage the semantic index to optimize pivot selection for some queries, rather than resorting to random pivot selection. This optimization is useful when there exists correlation between the rankings imposed by the user’s arbitrary sorting criteria and the rankings imposed by semantic similarity scores. In this case, LOTUS can sort the document set based on embedding distances to the user’s query, and select the -th item, rather than a random item, as the first pivot. This can reduce the number of LM comparisons required by subsequent rounds in the quick-select algorithm, leading to higher query efficiency at no accuracy loss. We believe fruitful future work will automatically estimate correlation between semantic similarity scores and the user-defined ranking criteria to transparently apply this optimization.
3.3. sem_join
Performing the semantic join involves evaluating the user’s natural language predicate on each pair of rows in the left and right table. Implemented naively, this can be prohibitively expensive, incurring a quadratic number of LM calls with respect to to the size of the left and right join tables. LOTUS thus implements several algorithms suitable for a variety of settings, highlighting the rich design space, which we plan to further optimize over in future work. Here we describe the design patterns currently implemented, which we find to be useful for real-world tasks that we evaluate in Section 4.
3.3.1. Algorithms
The first join algorithm implements the nested-loop join pattern with efficient LM batch processing to maximize GPU utilization rather than naively looping over each pair of rows and invoking the LM. This yields an LM call complexity, where and are the table sizes of the left and right join tables respectively. As Figure 18 shows, each LM call instructs the model to output a boolean value after evaluating the user’s natural language predicate. This quadratic join algorithm is suitable for small join tables.
Alternatively, we implement a map-search-filter join pattern which can be used to approximate the quadratic join algorithm, while incurring fewer LM calls. This algorithm first performs a semantic mapping over the left join key to the domain of the right join key. In the example provided by Figure 20, which joins paper abstracts with the datasets they use, the map step would invoke the LM over each abstract, instructing the model to output the dataset used. This step is un-grounded in the sense that the LM is not given knowledge of right join table, but may optionally leverage user-specified demonstrations. Following the semantic projection, the algorithm then leverages the semantic index and performs a similarity search over the right join table to find candidates likely to pass the predicate. In the example, this search would retrieve a list of datasets from the right table for each paper, using semantic similarity of the datasets to the LM projection output in the first step. The semantic similarity search is parameterized by , which is automatically set based on the user’s specified LM call budget. Lastly, the algorithm performs a filter over the candidate pairs and outputs the joined table of tuples pairs that pass. This join algorithm partially mimics the information flow of prior work (D’Oosterlinck et al., 2024) and abstracts away intermediate steps by instead allowing users to specify an LM call budget for the join. The LM call complexity of this algorithm is .
Lastly, we implement a search-filter join pattern. This algorithm first performs a semantic similarity join using a similarity index on the right join key, with the left join key embedded on the fly. In the example provided by Figure 20, this step would retrieve semantically similar datasets from the right table for each paper abstract in the left table. The batched semantic-similarity search sets the search parameter according to the user’s specified LM call budget, similar to the map-search-filter algorithm. The search results provide candidate tuple pairs, which are then evaluated in batch using the filter operation to gather the final set of rows that pass the user’s language predicate. This algorithm obtains an LM call complexity of .
3.3.2. Discussion
These join patterns offer performance tradeoffs suitable for different settings. The nested-loop algorithm offers an efficient solution when the join tables are sufficiently small, whereas the map-search-filter and search-filter patterns are suitable approximations that can apply efficiently over large tables, due to their linear LM call complexity in table size. The expected result quality of either approximation likely varies depending on the presence of predicate clustering and correlation, as defined by prior work (Patel et al., 2024) between the user’s predicate and the semantic embeddings. The search-filter pattern is likely to produce high quality results under positive query correlation, where entities that are semantically similar are also likely to pass the predicate. On the other hand, the map-search-filter pattern is likely to produce higher quality results when the predicate-embedding correlation is low using the original left join key’s embeddings as queries, but can be increased using the LM-projection over the left join key.
@@ 3.3.2. Discussion
Performing semantic aggregations are inherently challenging because, similar to the semantic top-, it requires logically reasoning across rows. Thus, the operator’s implementation must efficiently orchestrate the LM over large amounts of data, while managing long context inputs, which may degrade result quality (Liu et al., 2023) or overflow the underlying model’s context length. LOTUS aims to abstract away such low-level details from the user and provides an efficient implementation, designed to support high quality results and provide the programmer with flexibility.
3.4.1. Algorithms

### v2

@@ 3.2. Semantic Operators in LOTUS
Sem_topk & Sem_search . LOTUS supports a semantic top-k, which takes the langex ranking criteria, as described in Section 2. Programmers can optionally specify a group-by parameter to indicate a subset of columns to group over during ranking, as shown in Figure 4. The groupings are defined using standard equality matches over the group-by columns. Additionally, as the Figure shows, LOTUS also provides a top-k variant, sem_search, which assumes a semantic similarity-based ranking criteria relative to a natural language query. LOTUS also exposes advanced relevance-based re-ranking functionality for search, allowing users to specify the n_rerank parameter during the semantic search. The semantic search in this case will first find the top- most relevant documents and then re-rank the top- found documents to return the top n_rerank.
Sem_agg performs an aggregation over the input relation, with as langex signature that provides a commutative and associative aggregation function, as shown by Figure 2. Similar to sem_topk, LOTUS allows users to specify a group-by parameter.
Sem_group_by creates groups over the input dataframe according to the langex grouping criteria by clustering the data into a target number of groups (e.g., papers_df.sem_group_by("the topic of the abstract", 20)). By default, the operator automatically selects group labels, but the user can also optionally specify target labels. This operator is useful both for unsupervised discovery of semantic groups, and for semantic classification tasks.
Sem_map & Sem_extract both perform a natural language projection over an existing column. While sem_map projects to an arbitrary text attribute, sem_extract projects each tuple to a list of sub-strings from the source text. This is useful for applications, such as entity extraction or fact-checking, where finding snippets or verified quotes may be preferable to synthesized answers.
4. Execution Plans and Optimizations for Semantic Operators
Semantic operators create a rich design space of diverse execution plans which have significant consequences on system efficiency and accuracy. For example, there are many possible algorithms to join two collections of records using a LLM and a natural language expression (all the standard join algorithms, as well as ML heuristics to approximate the join, e.g., by semantically clustering similar records using an embedding model).
Specifically our goal is to offer an efficient plan while while guaranteeing high accuracy. Defining accuracy for semantic transformations, however, presents a fundamental challenge in the absence of ground truth labels. Therefore, we define accuracy of each operator plan with respect to a ”gold” algorithm, which uses a powerful oracle model . These ”gold” algorithms are tractable solutions that can achieve high quality results, but may be expensive in terms of execution time and LM call complexity. We discuss the design space of these ”gold” algorithms for several core semantic operators we introduced. In addition, we present several novel optimizations for semantic filters, joins, top-k and group-by. These optimizations are either lossless or provide statistical guarantees on accuracy with respect to a ”gold” algorithm with the expensive model . Specifically, our approximations emulate the results of the expensive but high-quality algorithm by using model cascades (Viola and Jones, 2001; Yue et al., 2024; Chen et al., 2023; Kang et al., 2017, 2022) or semantic indexes, building on past work on accelerating ML-based query processing with guarantees (Kang et al., 2022, 2019) but achieving this property for new types of operators. Notably, these optimizations can all be performed declaratively given a user’s accuracy targets. We show in Section 5 that they significantly accelerate real applications.
@@ 4. Execution Plans and Optimizations for Semantic Operators
While this paper proposes new optimizations for only a few, expensive semantic operators, we envision that a wide range of other new and traditional optimizations are possible under the semantic operator model. For example, several prior works demonstrate performance gains in both logical query plan optimizations (e.g. operator re-ordering (Liu et al., 2024c, b; Lin et al., 2024; Lu et al., 2018)) and other general LM-approximation techniques (e.g. code synthesis (Arora et al., 2023; Liu et al., 2024b) and prompt adaptation (Chen et al., 2023)).
@@
4.1.1. Gold Algorithm
Our gold algorithm for the semantic filter runs batched LLM calls over the tuples in relation . Each tuple, , is processed independently using the model , which is prompted to output a boolean value indicating whether the tuple passes the langex predicate.
4.1.2. Algorithmic Approximation
We can declaratively optimize semantic filters by providing an approximation which will obtain recall and precision targets and with respect to the gold implementation with probability . To achieve reduced cost we leverage the cheaper, but less accurate proxy model , which provides a score indicating whether the tuple passes the predicate, in conjunctions with the oracle LLM . This idea is inspired by prior works (Viola and Jones, 2001; Yue et al., 2024; Chen et al., 2023; Kang et al., 2017, 2022; Yue et al., 2024), which leverage model cascades for different problem settings. Specifically, several works study cascades in the video analytics setting with vision models, which have significantly different properties than LLMs, requiring different proxy scoring mechanisms. Additionally, other works, which study cascades for LLMs, use heavy-weight scoring mechanisms to decide whether to use the proxy model, which incurs a heavy execution time penalty, and does not provide accuracy guarantees. In contrast, we focus on providing accuracy guarantees for the constrained problem of applying cascades for filters, which allows us to use lighter-weight scoring functions than the more general case studied in prior works.
Providing accuracy guarantees is important because, in general, we cannot assume the proxy model is accurate. In fact, the proxy model, may perform very poorly for some tasks and datasets. The goal of our algorithm is to automatically discover the quality of the proxy by sampling and comparing to the gold algorithm with the oracle model. If the proxy provides high quality results, our algorithm will exploit it, leading to lower execution time. However, if the proxy model performs poorly, we will selectively route to it according to the proxy scores , and a learned threshold on the proxy scores. As prior work (Kang et al., 2020) discusses, such sampling-based algorithms to learn a threshold, , introduce multiple-hypothesis testing problems, requiring statistical corrections using confidence intervals, which we carefully apply in our algorithm.
Our algorithm leverages a small LLM as the proxy model. We generate scores using the the log-probabilities corresponding to the True or False output tokens of the proxy model’s answer and re-scaling by the quantiles over the proxy scores of tuples in the relation.
Algorithm 1 shows the full procedure for performing the approximate semantic filter. We begin by using an collecting a sample , which we label with the oracle and proxy models. The sampling procedure uses importance sampling and defensively mixes a uniform sample following prior work (Kang et al., 2022). Using a the central limit theorem and the normal approximation on the distribution of sample statistics, we then find thresholds and , which will respectively ensure the desired precision target and recall target are met, each with an error probability of . These sub-procedures follow the statistical techniques used in prior work (Kang et al., 2022), but applies them to this new setting where we must ensure both targets are met, requiring a correction for hypothesis testing and multiple failure modes.
Finally the algorithm proceeds to process each tuple in the relation. If the tuple’s proxy score is at least , we mark it as passing the predicate. If the tuple’s proxy score is less than or equal to , we mark it as failing the predicate, and for tuples with proxy scores between and , we resort to the oracle model to provide a label.
4.2. sem_join
4.2.1. Gold Algorithm
Performing the semantic join involves evaluating the user’s natural language predicate on each pair of rows in the left table, and right table, . The gold algorithm implements a nested-loop join pattern with efficient batched inference to maximize GPU utilization. This yields an LM call complexity, where and are the table sizes of the left and right join tables respectively. Each LM call instructs the model to output a boolean value after evaluating the user’s natural language predicate. This quadratic join algorithm is suitable for small join tables but scales poorly for large tables.
4.2.2. Algorithmic Approximation
Similar to semantic filters, we declaratively optimize semantic joins by providing an approximation that obtains recall and precision targets and with respect to the gold algorithm with probability . Due to the expensive quadratic scaling of the nested-loop join, rather than using a small LLM as the proxy model, we leverage an even cheaper proxy based on semantic similarity scores, which can be obtained efficiently with a semantic index over the join keys. We leverage two possible join approximation algorithms and dynamically chooses the lowest cost one, which will minimize the number of LLM calls.
The first approximation, sim-filter produces a proxy score , based on embedding similarity. We leverage the semantic index to perform batched similarity search and re-calibrate similarity scores according to the quantiles of the distribution over all pairs to obtain proxy scores. This approximation is likely to perform efficiently when tuple pairs with high semantic similarity between the the right and left join key are more likely to pass the predicate. This correlation phenomenon between predicate matches and embedding similarity has been studied by prior works (Patel et al., 2024) and is not always present. Thus, we introduce an alternative approximation suitable when correlation is not present.
The second approximation, map-sim-filter first performs a projection over the left join key and then uses the projected column to compute proxy scores uses the embedding similarity between the projected value and . As before we also re-calibrate the proxy scores taking their quantiles. Intuitively, performing the projection is useful when tuple pairs that pass the user’s natural language predicate exhibit low semantic similarity. The projection step invokes the LM over each tuple in left join table, prompting it to predict values in the attribute domain of the right join key that match. Notably, this LM projection is ungrounded, i.e., it is done without knowledge of the right join key’s attribute domain and can thus be performed in a fully parallelized semantic map operation, which can optionally take user-provided demonstrations. Figure 5 provides an example of a semantic join between a table of papers and datasets, where the predicate evaluates whether a given paper abstract uses a specific dataset. Here, the map step would invoke the LM over each abstract, instructing the model to output the dataset used, conditioned only on the abstract.
To dynamically choose between these two approximations, we follow a procedure similar to Algorithm 1 used for semantic filters. We begin by importance sampling to collect the sample and construct the set of oracle labels, , over the sample. We then obtain , and independently for each approximation algorithm using their respective proxy scores. Using the tuned thresholds and known proxy scores over for both algorithms over the dataset, we then determine the exact oracle cost needed to execute either algorithm, and we take the least cost plan with its associated learned thresholds to proceed to evaluate the predicate for each tuple pair.
4.3. sem_group_by
Performing this semantic operator entails discovering group labels and classifying each input tuple using the discovered labels. In general, performing the unsupervised group discovery is a clustering task, which is NP-hard (Dasgupta, [n. d.]). Our gold algorithm instead offers a tractable implementation to obtain group labels using a linear pass over the data with the reference model . Furthermore, once the labels are discovered, we perform the LM-based classification task and show how to optimize this step using an approximation with probabilistic guarantees.
4.3.1. Gold Algorithm
Our semantic group-by algorithm proceeds in two stages. The first stage performs the unsupervised label discovery, given the user’s grouping criteria and target number of groups. This stage first performs a semantic projection, prompting the LM to predict a label for each input tuple. Then, we create a semantic index over the projected column and perform a vector clustering over these predicted labels, using FAISS’ efficient k-means implementation to construct groups. Lastly, for each group, we top-k sample documents based on their similarity to the centroid, and few-shot prompt the LM to generate a label describing the given documents according to the user-provided grouping criteria (e.g. the topics of the documents) using a semantic aggregation.
In the second stage of the algorithm, we use the generated labels to perform classification over each document. We do this using a semantic map, which prompts the LM to output a label conditioned on the candidate labels and document.
4.3.2. Algorithmic Approximation
We provide an efficient approximation, specifically for the classification task in the second stage of the algorithm. Similar to semantic filters and join, we perform this optimization declaratively, while guaranteeing a classification accuracy target is met with probability . Similar, to semantic joins, we leverage semantic similarity scores as a cheap proxy. Specifically, we leverage the index constructed in stage one of the algorithm and compute similarity scores between the embeddings of each document’s un-grounded prediction and the embeddings of each class label. The proxy score of the tuple with a predicted label, , for group label is given by . Our goal is then to learn a threshold, , such that if the similarity between an embedding and the class label’s embedding is greater than , we return the class label, and otherwise we resort to the more expensive LM-based classification procedure. We accomplish this by uniform sampling and running the sub procedure equivalent to the used in Algorithm 1 for semantic filters, where the metric we evaluate is classification accuracy over the groups.
4.4. sem_topk
Performing a semantic top- ranking requires logically reasoning across rows, entailing joint reasoning over often large amounts of data. While prior works have studied LM-based passage re-ranking (Desai and Durrett, 2020; Drozdov et al., 2023; Liang et al., 2022; Ma et al., 2023; Pradeep et al., 2023a, b; Qin et al., 2024; Sachan et al., 2022; Sun et al., 2023) and ranking with noisy comparisons (Shah and Wainwright, 2016; Braverman and Mossel, [n. d.]) with the goal of achieving high quality results in a modest number of total LM calls or comparisons, our implementations provides a generalized ranking algorithms for arbitrary natural language expressions and aims to produce both high quality results and low execution time. Here we outline leverage several algorithmic design decisions for the gold algorithm as well as a lossless optimization.
4.4.1. Gold Algorithm
Two important algorithmic design decisions for the semantic top-k include how to implement the LM-based comparison and how to aggregate ranking information across LM-based comparisons. Our implementation uses pairwise LM-based comparisons for the former, and a quick-select based top-k algorithm (Hoare, 1961) for the latter.
We briefly describe the reason for these choices and alternatives considered. Pairwise-prompting methods offer a simple and effective approach that feeds a single pair of documents to the LM in each invocation, prompting the model to perform a comparison and output a binary label. The two main classes of alternatives are point-wise ranking methods (Desai and Durrett, 2020; Drozdov et al., 2023; Liang et al., 2022; Sachan et al., 2022; Wu et al., 2024), and list-wise ranking methods (Ma et al., 2023; Pradeep et al., 2023a, b; Sun et al., 2023), both of which have been shown to face quality issues (Sun et al., 2023; Qin et al., 2024; Desai and Durrett, 2020). In contrast, pairwise comparisons have been shown to be an effective base unit for ranking with relatively high robustness to input order sensitivity (Qin et al., 2024).
In addition, we studied several alternative rank-aggregation algorithm, including quadratic sorting algorithms, a heap-based top-k algorithm and a quick-select-based top-k ranking algorithm, which we discuss in Section 5. The quick-select based algorithm proceeds in successive rounds, each time choosing a pivot, and comparing all other remaining elements in the document set to the pivot item to determine the rank of the pivot. Because each round is fully parallelizable, we can efficiently batch these LM-based comparisons before recursing in the next round. In section 5, we find that the quick-select-based algorithm offers high accuracy while also offering an efficient implementation with at least an order magnitude fewer total calls then the quadratic sorting algorithm and more opportunities for efficient batched inference, leading to lower execution time, compared to a heap-based implementation.
@@ 4.4.2. Optimizations
We leverage the semantic index to optimize pivot selection for some queries, while incurring no accuracy loss. This optimization is useful when there exists correlation between the rankings imposed by the user’s arbitrary sorting criteria and the rankings imposed by semantic similarity scores. In this case, LOTUS can sort the document set based on embedding distances to the user’s query, and select the -th item, rather than a random item, as the first pivot. This can reduce the number of LM comparisons required by subsequent rounds in the quick-select algorithm, leading to higher query efficiency at no accuracy loss. In the case of no correlation between the langex-based ranking and similarity-based ranking, this method amounts to random pivot selection, and in the worst case of an adversarial pivot, the algorithm will incur one extra pivot round, which can impact execution time, but not degrade quality.
@@
Performing semantic aggregations is inherently challenging because, similar to the semantic top-, it requires logically reasoning across rows. Thus, the operator’s implementation must efficiently orchestrate the LM over large amounts of data, while managing long context inputs, which may degrade result quality (Liu et al., 2023) or overflow the underlying model’s context length. LOTUS aims to abstract away such low-level details from the user and provides an efficient implementation, designed to support high quality results.
4.5.1. Gold Algorithm